Brock Florist Company buys a new delivery truck for ​$34,000. It is classified as a​ light-duty truck. the tax rate is 0.3

a. Calculate the depreciation schedule using a​ five-year life and MACRS depreciation.

b. Calculate the remaining book value at the end of year 4 and the termination value​ (aka. cost recovery or after tax salvage​ value) for the delivery truck if the truck can be sold at the end of year 4 ​for$2,000.

a.  Calculate the depreciation schedule using a​ five-year life and MACRS depreciation.​ (Round to the nearest​ dollar.)

Annual Depreciation

Year 1

​$

Year 2

​$

Year 3

​$

Year 4

​$

Year 5

​$

Year 6

​$

b. Calculate the remaining book value at the end of year 4 and the termination value​ (aka. cost recovery or after tax salvage​ value) at the end of year 4 for the delivery truck.

Remaining book value at the end of year 4 is ​$5,875​(Round to the nearest​ dollar.)The termination value​ (cost recovery or after tax salvage​ value) at the endo of year 4 is ​$3,163  ​(Round to the nearest​ dollar.)
Answer #1
Year Cost of Asset Applicable Depreciation % Depreciation Amount Remaining Book value
1 34000 20% 6800 27200
2 34000 32% 10880 16320
3 34000 19.20% 6528 9792
4 34000 11.52% 3916.8 5875.2
5 34000 11.52% 3916.8 1958.4
6 34000 5.76% 1958.4 0
Remaininhg book value after year 4 = $5875.2
Selling Price = 2000
Loss on Sale = $3875.2
Savings of tax on loss = 3875.2*0.3 = $1162.56
After tax salvage value = 2000+1162.56 = 3162.56
i.e. $3163